# Ferret: Federated Full-Parameter Tuning at Scale for Large Language Models

### v1

@@ 3.2 Update Projection and Reconstruction
Choice of Random Bases ${\mathbf{V}}$. Particularly, if ${\mathbf{V}}$ is a rectangular matrix with ones on its main diagonal, meaning that each ${\mathbf{v}}_{k}$ is a standard basis vector, (5) simplifies to ${\bm{\gamma}}={\mathbf{V}}^{\top}\Delta$, which then corresponds to a block-wise dimension selection for local update projection and reconstruction. However, this approach significantly reduces the number of parameters updated per round as $K\ll d$, potentially hindering the overall tuning performance. We thus propose to sample each element in ${\mathbf{v}}_{k}\ (k\in[K])$ independently from a normal distribution with bounded 2-norm, i.e., $\left\|{\mathbf{v}}_{k}\right\|\leq 1$, aiming to realize and stabilize full-parameter tuning of LLMs for competitive overall performance. To achieve this, we can sample from a truncated normal distribution: ${\textnormal{v}}\sim{\mathcal{N}}(0,1)$ with ${\textnormal{v}}\in[-1/\sqrt{d},1/\sqrt{d}]$ instead. The efficacy of this bounded norm will be demonstrated in Sec. 4.1.
Reconstruction w/o Inversion. Unfortunately, (5) incurs a computational complexity of ${\mathcal{O}}(K^{2}d+K^{3})$ and storage complexity of ${\mathcal{O}}(Kd)$ owing to the inversion of ${\mathbf{V}}^{\top}{\mathbf{V}}$ in (5), which is prohibitively costly, especially when $K$ is large and $d$ reaches billions. Since ${\mathbf{V}}^{\top}{\mathbf{V}}$ is a scaled empirical covariance for the aforementioned distribution of an identity covariance matrix [19] , we propose to approximate ${\mathbf{V}}^{\top}{\mathbf{V}}$ with ${\mathbf{I}}_{K}$ (i.e., $K\times K$-dimensional identity matrix) and (5) as
@@
Here, $\rho_{l}\triangleq 1-\frac{2\psi(1/\sqrt{d_{l}})/\sqrt{d_{l}}}{2\Phi(1/\sqrt{d_{l}})-1}$. This trick reduces the storage complexity to ${\mathcal{O}}(\max\{\{K_{l},d_{l}\}_{l=1}^{\smash{L}}\})$ that is straightforward to verify, and lowers the computational complexity to ${\mathcal{O}}(\sum_{l\in[L]}K_{l}d_{l})$. Of note, (8) also significantly reduces the computational complexity of global aggregation compared to existing methods [12] (verified in Sec. 5). This block-wise reconstruction thus further enhances the scalability of our Ferret in the federated full-parameter tuning of LLMs.
@@ 4 Theoretical Analyses and Insights
We then provide theoretical analyses to substantiate the effectiveness of Ferret: (a) reconstruction analysis in Sec. 4.1; (b) convergence analysis in Sec. 4.2; and (c) scalability and beyond in Sec. 4.3.
@@ 4.1 Reconstruction Analysis
Given the reconstruction in (7), we have
$\mathbb{E}\left[\widetilde{\Delta}\right]=\Delta\ .$
To begin with, we demonstrate in Thm. 1 that our reconstruction in (7) is unbiased, with the proof provided in Appx. B.1. Of note, Thm. 1 shows that (a) the scalar $1/(\rho K)$ is crucial for (7) to achieve an unbiased reconstruction of the ground-truth update $\Delta$, and (b) our (7) avoids the bias commonly found in zeroth-order FL methods [14], including FedZO [20] and FedKSeed [12]. As a result, (7) is expected to provide a more accurate update reconstruction, which we will elaborate more below.
@@
Given the reconstruction in (7), we have
$\mathbb{E}\left[\left\|\widetilde{\Delta}-\Delta\right\|\right]\leq\max\left\{2\sqrt{\frac{2\ln(2d)}{\rho K}},\frac{2\ln(2d)}{\rho K}\right\}\left\|\Delta\right\|\ .$
We then demonstrate the efficacy of our reconstruction in (7) by theoretically bounding the difference between the reconstructed update $\widetilde{\Delta}$ and the ground truth $\Delta$ in Thm. 2. The proof is in Appx. B.2. Of note, $1/\rho$ typically has an asymptotic rate of ${\mathcal{O}}(d)$, which we will verify empirically in Appx. C.4. Thm. 2 offers three critical insights of our Ferret: (a) Our reconstruction in (7) incurs a reconstruction error at a rate of $\widetilde{{\mathcal{O}}}(d/K)$ for $T$ local update iterations when $\sqrt{d}\geq K$, which generally aligns with the results in [21]. This indicates that the reconstruction error of our (7) can be linearly reduced by increasing $K$. (b) Ferret avoids additional constant error items [14] that are caused by the biased estimation in these zeroth-order FL methods, implying that our (7) can be more accurate. We will justify this further in our Thm. 3 below. (c) Thanks to the independence from the iterations (i.e., $T$) of local updates in Thm. 2, Ferret prevents the error accumulation over the local update iterations $T$, which is a common issue in zeroth-order FL methods [20, 12].
@@
Define $g_{k}\triangleq\frac{\ell({\mathbf{w}}+\epsilon{\mathbf{v}}_{k};{\mathbf{x}}^{(i)})-\ell({\mathbf{w}};{\mathbf{x}}^{(i)})}{\epsilon}$ where each element v in ${\mathbf{v}}_{k}$ is sampled from ${\textnormal{v}}\sim{\mathcal{N}}(0,1)$ with ${\textnormal{v}}\in[-1/\sqrt{d},1/\sqrt{d}]$, ${\bm{g}}\triangleq[g_{1}\ \cdots\ g_{K}]^{\top}$, and ${\mathbf{V}}\triangleq[{\mathbf{v}}_{1}\ {\mathbf{v}}_{2}\ \cdots\ {\mathbf{v}}_{K}]\in{\mathbb{R}}^{d\times K}$, assume $\ell(\cdot;\cdot)$ is $\beta$-smooth w.r.t its first argument, the zeroth-order reconstruction ${\mathbf{V}}{\bm{g}}/K$ used in [20, 12] then incurs:
$\left\|\frac{1}{K}{\mathbf{V}}{\bm{g}}-\frac{1}{K}{\mathbf{V}}{\mathbf{V}}^{\top}\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})\right\|\leq\frac{1}{2}\beta\epsilon\ .$
We then show in Thm. 3 the connection between our update projection (6) and zeroth-order method used in [20, 12]. The proof is provided in Appx. B.3. Thm. 3 delivers three essential insights: (a) When ${\epsilon}\rightarrow 0$, the reconstruction ${\mathbf{V}}{\bm{g}}/K$ in zeroth-order method is equivalent to ${\mathbf{V}}{\mathbf{V}}^{\top}\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})/K$ and shares a similar form of (7) when $\Delta$ is replaced by $\nabla\ell({\mathbf{w}};{\mathbf{x}}^{\smash{(i)}})$, implying that zeroth-order method in fact aims to approximate our reconstruction (7). (b) In practice, $\epsilon>0$. So, zeroth-order method leads to a biased reconstruction with an additional error term of $\beta{\epsilon}/2$ compared to our (7), and this error will accumulate over $T$ local iterations, implying that our (7) can indeed be more accurate as we have demonstrated above. (c) In addition, zeroth-order method is typically coupled with a single gradient (i.e., $\nabla\ell({\mathbf{w}};{\mathbf{x}}^{\smash{(i)}})$), whereas our (7) can be applied to any vector, making it more general. Overall, these results further verify the advantages of our (7) over the zeroth-order method used in [20, 12], which we will also support empirically in Appx. C.4.
@@
For block-wise reconstruction (8) of size $L$,
$\sum_{l\in[L]}d_{l}K_{l}<\Bigg{(}\sum_{l\in[L]}d_{l}\Bigg{)}\Bigg{(}\sum_{l\in[L]}K_{l}\Bigg{)}=dK\ .$
@@
For block-wise reconstruction (8) of size $L$, when $\sqrt{d_{l}}\geq K_{l}$ for any $l\in[L]$,
$\mathbb{E}\left[\left\|\widetilde{\Delta}-\Delta\right\|\right]<\widetilde{{\mathcal{O}}}\Bigg{(}\sum_{l\in[L]}\frac{\left\|\Delta_{l}\right\|}{\rho_{l}K_{l}}\Bigg{)}\ ,$
which is minimized by choosing $K_{l}\propto\sqrt{\left\|\Delta_{l}\right\|/\rho_{l}}$.
@@ 4.2 Convergence Analysis
Define $D\triangleq{\mathcal{L}}({\mathbf{w}}_{0})-\min_{{\mathbf{w}}}{\mathcal{L}}({\mathbf{w}})$. Assume that ${\mathcal{L}}({\mathbf{w}})$ is $\beta$-smooth and non-convex, and $\mathbb{E}[\|\nabla{\mathcal{L}}^{\smash{(i)}}({\mathbf{w}})-\ell({\mathbf{w}};{\mathbf{x}})\|^{2}]\leq\sigma^{2}$ for any ${\mathbf{x}},{\mathbf{w}}$, when choosing $\eta\leq\frac{1}{20\beta T}$ in Algo. 1, the following holds for federated full-parameter tuning with ${\mathcal{L}}^{\smash{(i)}}({\mathbf{w}})={\mathcal{L}}({\mathbf{w}})$,
$\min_{r\in[R)}\mathbb{E}\left[\left\|\nabla{\mathcal{L}}({\mathbf{w}}_{r})\right\|^{2}\right]\leq{\mathcal{O}}\left(\frac{D}{\eta TR}+\eta T\sigma^{2}\right)\ .$
Especially, by choosing $\eta=\frac{1}{20\beta T\sqrt{R}}$ in Algo. 1, the number of communication rounds are required to be $R={\mathcal{O}}(1/{\epsilon}^{2})$ to achieve an ${\epsilon}$ convergence error.
Its proof is in Appx. B.6. Particularly, when $T=1$, Thm. 4 recovers the result of standard SGD [22]. Thm. 4 provides three essential insights: (a) Thanks to our improved update reconstruction (7) as justified above, Ferret avoids the additional constant terms accumulated over $T$ local iterations, which are typically caused by the biased gradient estimation in zeroth-order FL methods (e.g., FedZO and FedKSeed) [20], thereby highlighting the superior advantage of Ferret over these zeroth-order FL methods in convergence speed. (b) Given a proper $\eta$, Ferret shares the same communication round complexity as SGD, at a rate of ${\mathcal{O}}(1/\epsilon^{\smash{2}})$, showing that the communication round complexity of Ferret is asymptotically comparable to that of standard SGD. (c) This communication rounds complexity is improved over that of zeroth-order FL methods [20] due to its independence from $d$ and other constant factors required by these zeroth-order FL methods, further highlighting the advantage of Ferret in communication round complexity and its improved efficacy in federated full-parameter tuning over these methods.
4.3 Scalability and Beyond
Table 1: Comparison of scalability (computation and communication per round, and ##\##rounds to converge) and other factors (adaptivity, generalization, and privacy). Here, $d\gg K\gg T$. Symbols: $\circ$ (fewer is better), $\heartsuit$ and $\diamond$ (more is better).
@@ 4.3 Scalability and Beyond
Rounds to Converge. As revealed in Sec.4.2, our Ferret benefits from unbiased update reconstruction in (7) (validated in Thm. 1), enabling fast convergence with a small number of communication rounds to achieve $\epsilon$ convergence error (see Thm. 4). This is significantly more efficient than zeroth-order FL methods like FedZO and FedKSeed, which require many more communication rounds to converge due to poor gradient estimation [20]. FedAvg, applying the ground truth local update for its global aggregation, surely converges with the fewest rounds. Overall, Ferret remains a strong choice for federated full-parameter tuning of LLMs, even in terms of communication rounds to converge.
Beyond Scalability. Our Ferret also offers benefits in adaptability, generalization, and privacy. Unlike FedKSeed, which is limited to SGD, Ferret is highly adaptable, because both global aggregation (2) and local update (3) in Ferret can be implemented with any gradient method variant, e.g., the widely used AdamW [23] in LLM training. This adaptability thus makes it much easier to integrate Ferret into existing centralized tuning workflows for LLMs, facilitating a seamless transition to federated tuning. Besides, since Ferret enables federated tuning with full parameters, it is expected to deliver strong generalization performance as other federated full-parameter tuning methods like FedAvg, as supported in Sec. 5. Finally, by transmitting only seeds and low-dimensional projected coordinates among clients, rather than the entire model (or gradients) as in FedZO and FedAvg, Ferret ensures improved privacy, like FedKSeed, for federated full-parameter tuning of LLMs.
Overall, Ferret strikes an optimal balance between computational efficiency, communication overhead, convergence speed, and other critical factors such as adaptability, generalization, and privacy. This makes it a highly scalable and desirable solution for federated full-parameter tuning of LLMs.

### v3

@@ 3.2 Update Projection and Reconstruction
Choice of Random Bases ${\mathbf{V}}$. Particularly, if ${\mathbf{V}}$ is a rectangular matrix with ones on its main diagonal, meaning that each ${\mathbf{v}}_{k}$ is a standard basis vector, (5) simplifies to ${\bm{\gamma}}={\mathbf{V}}^{\top}\Delta$, which then corresponds to a block-wise dimension selection for local update projection and reconstruction. However, this approach significantly reduces the number of parameters updated per round as $K\ll d$, potentially hindering the overall tuning performance. We thus propose to sample each element in ${\mathbf{v}}_{k}\ (k\in[K])$ independently from a normal distribution with bounded 2-norm, i.e., $\left\|{\mathbf{v}}_{k}\right\|\leq 1$, aiming to realize and stabilize full-parameter tuning of LLMs for competitive overall performance. To achieve this, we can sample from a truncated normal distribution: ${\textnormal{v}}\sim{\mathcal{N}}(0,1)$ with ${\textnormal{v}}\in[-1/\sqrt{d},1/\sqrt{d}]$ instead. The efficacy of this bounded norm will be demonstrated in Sec. 4.1 shortly.
Reconstruction w/o Inversion. Unfortunately, (5) incurs a computational complexity of ${\mathcal{O}}(K^{2}d+K^{3})$ and storage complexity of ${\mathcal{O}}(Kd)$ owing to the inversion of ${\mathbf{V}}^{\top}{\mathbf{V}}$ in (5), which is prohibitively costly, especially when $K$ is large and $d$ reaches billions. Since ${\mathbf{V}}^{\top}{\mathbf{V}}$ is a scaled empirical covariance for the aforementioned distribution of an identity covariance matrix (Vershynin, 2012) , we propose to approximate ${\mathbf{V}}^{\top}{\mathbf{V}}$ with ${\mathbf{I}}_{K}$ (i.e., $K\times K$-dimensional identity matrix) and (5) as
@@
Here, $\rho_{l}\triangleq 1-\frac{2\psi(1/\sqrt{d_{l}})/\sqrt{d_{l}}}{2\Phi(1/\sqrt{d_{l}})-1}$. This trick reduces the storage complexity to ${\mathcal{O}}(\max\{\{K_{l},d_{l}\}_{l=1}^{\smash{L}}\})$ that is straightforward to verify, and lowers the computational complexity to ${\mathcal{O}}(\sum_{l\in[L]}K_{l}d_{l})$. Of note, (8) also significantly reduces the computational complexity of global aggregation compared to existing methods (Qin et al., 2024) (verified in Sec. 5). This block-wise reconstruction thus further enhances the scalability of our Ferret in the federated full-parameter tuning of LLMs. Interestingly, this block-wise strategy has also been widely applied in other fields due to it efficacy (Frantar & Alistarh, 2023; Luo et al., 2024).
@@ 4 Theoretical Analyses and Insights
We now provide theoretical analyses to substantiate the effectiveness of Ferret: (a) reconstruction analysis in Sec. 4.1; (b) convergence analysis in Sec. 4.2; and (c) scalability and beyond in Sec. 4.3.
@@ 4.1 Reconstruction Analysis
Given the reconstruction in (7), we have $\mathbb{E}\left[\widetilde{\Delta}\right]=\Delta\ .$
To begin with, we demonstrate in Thm. 1 that our reconstruction in (7) is unbiased, with the proof provided in Appx. B.1. Of note, Thm. 1 shows that (a) the scalar $1/(\rho K)$ is crucial for (7) to achieve an unbiased reconstruction of the ground-truth update $\Delta$, and (b) our (7) avoids the bias commonly found in zeroth-order FL methods (Berahas et al., 2022), including FedZO (Fang et al., 2022) and FedKSeed (Qin et al., 2024). As a result, (7) is expected to provide a more accurate update reconstruction as shown below.
@@
Given the reconstruction in (7), we have $\mathbb{E}\left[\left\|\widetilde{\Delta}-\Delta\right\|\right]\leq\max\left\{2\sqrt{\frac{2\ln(2d)}{\rho K}},\frac{2\ln(2d)}{\rho K}\right\}\left\|\Delta\right\|\ .$
We then demonstrate the efficacy of our reconstruction in (7) by theoretically bounding the difference between the reconstructed update $\widetilde{\Delta}$ and the ground truth $\Delta$ in Thm. 2. The proof is in Appx. B.2. Of note, $1/\rho$ typically has an asymptotic rate of ${\mathcal{O}}(d)$, which we will verify empirically in Appx. C.5. Thm. 2 offers three critical insights of our Ferret: (a) Our reconstruction in (7) incurs a reconstruction error at a rate of $\widetilde{{\mathcal{O}}}(d/K)$ for $T$ local update iterations when $\sqrt{d}\geq K$, which generally aligns with the results in (Vershynin, 2010). This indicates that the reconstruction error of our (7) can be linearly reduced by increasing $K$. (b) Ferret avoids additional constant error items (Berahas et al., 2022) that are caused by the biased estimation in these zeroth-order FL methods, implying that our (7) can be more accurate. We will justify this further in our Thm. 3 below. (c) Thanks to the independence from the iterations (i.e., $T$) of local updates in Thm. 2, Ferret prevents the error accumulation over the local update iterations $T$, which is a common issue in zeroth-order FL methods (Fang et al., 2022; Qin et al., 2024).
@@
Define $g_{k}\triangleq\frac{\ell({\mathbf{w}}+\epsilon{\mathbf{v}}_{k};{\mathbf{x}}^{(i)})-\ell({\mathbf{w}};{\mathbf{x}}^{(i)})}{\epsilon}$, in which each element v in ${\mathbf{v}}_{k}$ is sampled from ${\textnormal{v}}\sim{\mathcal{N}}(0,1)$ with ${\textnormal{v}}\in[-1/\sqrt{d},1/\sqrt{d}]$, ${\bm{g}}\triangleq[g_{1}\ \cdots\ g_{K}]^{\top}$, and ${\mathbf{V}}\triangleq[{\mathbf{v}}_{1}\ {\mathbf{v}}_{2}\ \cdots\ {\mathbf{v}}_{K}]\in{\mathbb{R}}^{d\times K}$, assume $\ell(\cdot;\cdot)$ is $\beta$-smooth w.r.t its first argument, the zeroth-order reconstruction ${\mathbf{V}}{\bm{g}}/K$ used in (Fang et al., 2022; Qin et al., 2024) then incurs: $\left\|\frac{1}{K}{\mathbf{V}}{\bm{g}}-\frac{1}{K}{\mathbf{V}}{\mathbf{V}}^{\top}\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})\right\|\leq\frac{1}{2}\beta\epsilon\ .$
We then show in Thm. 3 the connection between our update projection (6) and zeroth-order method used in (Fang et al., 2022; Qin et al., 2024). The proof is provided in Appx. B.3. Thm. 3 delivers three essential insights: (a) When ${\epsilon}\rightarrow 0$, the reconstruction ${\mathbf{V}}{\bm{g}}/K$ in zeroth-order method is equivalent to ${\mathbf{V}}{\mathbf{V}}^{\top}\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})/K$ and shares a similar form of (7) when $\Delta$ is replaced by $\nabla\ell({\mathbf{w}};{\mathbf{x}}^{\smash{(i)}})$, implying that zeroth-order method in fact aims to approximate our reconstruction (7). (b) In practice, $\epsilon>0$. So, zeroth-order method leads to a biased reconstruction with an additional error term of $\beta{\epsilon}/2$ compared to our (7), and this error will accumulate over $T$ local iterations, implying that our (7) can indeed be more accurate as we have demonstrated above. (c) In addition, zeroth-order method is typically coupled with a single gradient (i.e., $\nabla\ell({\mathbf{w}};{\mathbf{x}}^{\smash{(i)}})$), whereas our (7) can be applied to any vector, making it more general. Overall, these results further verify the advantages of our (7) over the zeroth-order method used in (Fang et al., 2022; Qin et al., 2024), which we will also support empirically in Appx. C.5.
@@
For block-wise reconstruction (8) of size $L$, $\sum_{l\in[L]}d_{l}K_{l}<\Bigg{(}\sum_{l\in[L]}d_{l}\Bigg{)}\Bigg{(}\sum_{l\in[L]}K_{l}\Bigg{)}=dK\ .$
@@
For block-wise reconstruction (8) of size $L$, when $\sqrt{d_{l}}\geq K_{l}$ for any $l\in[L]$, $\mathbb{E}\left[\left\|\widetilde{\Delta}-\Delta\right\|\right]<\widetilde{{\mathcal{O}}}\Bigg{(}\sum_{l\in[L]}\frac{\left\|\Delta_{l}\right\|}{\rho_{l}K_{l}}\Bigg{)}\ ,$ which is minimized by choosing $K_{l}\propto\sqrt{\left\|\Delta_{l}\right\|/\rho_{l}}$.
@@ 4.2 Convergence Analysis
Define $D\triangleq{\mathcal{L}}({\mathbf{w}}_{0})-\min_{{\mathbf{w}}}{\mathcal{L}}({\mathbf{w}})$. Assume that ${\mathcal{L}}({\mathbf{w}})$ is $\beta$-smooth and non-convex, and $\mathbb{E}[\|\nabla{\mathcal{L}}^{\smash{(i)}}({\mathbf{w}})-\nabla\ell({\mathbf{w}};{\mathbf{x}})\|^{2}]\leq\sigma^{2}$ for any ${\mathbf{x}},{\mathbf{w}}$, when choosing $\eta\leq\frac{1}{20\beta T}$ in Algo. 1, the following holds for federated full-parameter tuning with ${\mathcal{L}}^{\smash{(i)}}({\mathbf{w}})={\mathcal{L}}({\mathbf{w}})$, $\min_{r\in[R)}\mathbb{E}\left[\left\|\nabla{\mathcal{L}}({\mathbf{w}}_{r})\right\|^{2}\right]\leq{\mathcal{O}}\left(\frac{D}{\eta TR}+\eta T\sigma^{2}\right)$ where $[R)$ is the half-open interval $[0,R)$. Especially, by choosing $\eta=\frac{1}{20\beta T\sqrt{R}}$ in Algo. 1, the number of communication rounds are required to be $R={\mathcal{O}}(1/{\epsilon}^{2})$ to achieve an ${\epsilon}$ convergence error.
Its proof is in Appx. B.6. Particularly, when $T=1$, Thm. 4 recovers the result of standard SGD (Ghadimi & Lan, 2013). Thm. 4 provides three essential insights: (a) Thanks to our improved update reconstruction (7) as justified above, Ferret avoids the additional constant terms accumulated over $T$ local iterations, which are typically caused by the biased gradient estimation in zeroth-order FL methods (e.g., FedZO and FedKSeed) (Fang et al., 2022), thereby highlighting the superior advantage of Ferret over these zeroth-order FL methods in convergence speed. (b) Given a proper $\eta$, Ferret shares the same communication round complexity as SGD, at a rate of ${\mathcal{O}}(1/\epsilon^{\smash{2}})$, showing that the communication round complexity of Ferret is asymptotically comparable to that of standard SGD. (c) This communication rounds complexity is improved over that of zeroth-order FL methods (Fang et al., 2022) due to its independence from $d$ and other constant factors required by these zeroth-order FL methods, further highlighting the advantage of Ferret in communication round complexity and its improved efficacy in federated full-parameter tuning over these methods.
Table 1: Comparison of scalability (computation and communication per round, and ##\##rounds to converge) and other factors (adaptability, generalization, and privacy). Here, $d\gg K\gg T$. Symbols: $\circ$ (fewer is better), $\heartsuit$ and $\diamond$ (more is better).
@@ 4.3 Scalability and Beyond
Rounds to Converge. As revealed in Sec.4.2, our Ferret benefits from unbiased update reconstruction in (7) (validated in Thm. 1), enabling fast convergence with a small number of communication rounds to achieve $\epsilon$ convergence error (see Thm. 4). This is significantly more efficient than zeroth-order FL methods like FedZO and FedKSeed, which require many more communication rounds to converge due to poor gradient estimation (Fang et al., 2022). FedAvg, applying the ground truth local update for its global aggregation, surely converges with the fewest rounds. Overall, Ferret remains a strong choice for federated full-parameter tuning of LLMs, even in terms of rounds to converge.
Beyond Scalability. Our Ferret also offers benefits in adaptability, generalization, and privacy. Unlike FedKSeed, which is limited to SGD, Ferret is highly adaptable, because both global aggregation (2) and local update (3) in Ferret can be implemented with any gradient method variant, e.g., the widely used AdamW (Loshchilov & Hutter, 2019) in LLM training. This adaptability thus makes it much easier to integrate Ferret into existing centralized tuning workflows for LLMs, facilitating a seamless transition to federated tuning. Besides, since Ferret enables federated tuning with full parameters, it is expected to deliver strong generalization performance as other federated full-parameter tuning methods like FedAvg, as supported in Sec. 5. Finally, by transmitting only seeds and low-dimensional projected coordinates among clients, rather than the entire model (or gradients) as in FedZO and FedAvg, Ferret ensures improved privacy for federated full-parameter tuning of LLMs.
Overall, Ferret strikes an optimal balance between computational efficiency, communication overhead, convergence speed, and other critical factors such as adaptability, generalization, and privacy. This makes it a scalable and desirable solution for federated full-parameter tuning of LLMs.